# Asymptotic geometry of non-abelian Hodge theory and Riemann–Hilbert correspondence, rank three $\widetilde{E}_{6}$ case

### v1

@@ 2.3 The Hitchin fibration
According to Lemma 2.5, with the choices made in Section 2.2, the first two components vanish. The third characteristic coefficient of $\theta$ is $\operatorname{det}\theta\in H^{0}(\mathbb{C}P^{1},K(D)^{\otimes 3})$. Let us use the notation $L=K(D)$, with the natural projection from the total space of $L$, $p_{L}:\textrm{Tot}L\rightarrow\mathbb{C}P^{1}$. Define $\zeta\frac{\operatorname{d}\!z}{z(z-1)}$ to be the canonical section of $p_{L}^{*}L$ over $p_{L}^{-1}(\mathbb{C})$, i.e. away from the infinity section. With this, the characteristic polynomial of $\theta$ is
@@
where $\operatorname{det}\theta$ may take its values is called the Hitchin base of $\mathcal{M}_{\textrm{Dol}}(\alpha)$, and the map
$H:\mathcal{M}_{\textrm{Dol}}(\alpha)\rightarrow\mathcal{H},\hskip 8.5359pt(\mathcal{E},\theta)\mapsto\operatorname{det}\theta$
@@
$\Sigma_{t}=\{(z,\zeta)|\zeta^{3}+tz(z-1)=0\}\subset\textrm{Tot}L.$ (8)
Notice that $\Sigma_{t}$ has maximal ramification over $0,1$ and a smooth compactification, denoted by $\widetilde{\Sigma}_{t}$, in $\textrm{Tot}L$ at $z=\infty$ with $(w,\zeta)=(0,0)$, where it also admits cyclic ramification. One can easily compute from Riemann–Hurwitz formula that its genus is equal to 1.
@@ 2.4 Ramification of the spectral curve and the Jacobian
$\textrm{deg}\mathcal{L}=\textrm{deg}\mathcal{E}+r(1-r)\left(1-g-\frac{n}{2}\right)=3,$
where $r=3$ is the rank, $n=3$ is the number of ramification points, and $g=0$ is the genus of $\mathbb{C}P^{1}$. Thus an element of the Dolbeault moduli space determines a spectral curve and a line bundle of degree 3 on it, and vice versa. So, the fiber of the Hitchin fibration over the point parameterized by a fixed $t\in\mathcal{H}$ is $\textrm{Pic}^{3}(\widetilde{\Sigma}_{t})$, that is a 2-torus, namely a torsor over $\operatorname{Jac}(\widetilde{\Sigma}_{t})$. In particular, the Hitchin fibration is an elliptic fibration.
Let us discuss this correspondence between the Hitchin fibers and the Jacobian of the spectral curve a bit more in detail. Consider the following period lattice $\Lambda_{t}\subset H^{0,1}(\widetilde{\Sigma}_{t})\cong\mathbb{C}$, provided by the image $\textrm{Im}(p^{0,1}\circ\iota)$, where
@@
are the coefficient inclusion on the first cohomology class, and the projection of harmonic forms to their antiholomorphic part respectively. There exists a $\mathbb{C}$-analytic isomorphism $\operatorname{Jac}(\widetilde{\Sigma}_{t})\cong H^{0,1}(\widetilde{\Sigma}_{t})/\Lambda_{t}$. Namely, any class in $H^{0,1}(\widetilde{\Sigma}_{t})/\Lambda_{t}$ can be represented by a $\mu\in\Omega^{0,1}(\widetilde{\Sigma}_{t})$, because of the abelian Hodge correspondence. Then the connection form $u=\mu-\overline{\mu}=2i\textrm{Im}\mu\in\Omega^{1}(\widetilde{\Sigma}_{t})$ defines a flat $U(1)$-connection. For the line bundle $\mathcal{L}\in\operatorname{Jac}(\widetilde{\Sigma}_{t})$ given by $\mu$ and a $1$-cycle $X\in H_{1}(\widetilde{\Sigma}_{t},\mathbb{Z})$ we call
@@ 2.5 The Hitchin section
$\textrm{res}_{P}(\tau\theta_{1})=\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&1&0\end{pmatrix}.$
@@ 3 Description of the Betti moduli space
In this chapter we consider the Betti space (or character variety), parameterizing the irreducible representations of the fundamental group of $\mathbb{C}P^{1}\setminus D$ in $\textrm{GL}(3,\mathbb{C})$, up to the simultaneous conjugation by of $\textrm{PGL}(3,\mathbb{C})$. We pick $z_{0}\in\mathbb{C}P^{1}\setminus\{0,1,\infty\}$ once and for all, and all occurrences of fundamental group will mean with base point $z_{0}$. Since the fundamental group of $\mathbb{C}P^{1}\setminus\{0,1,\infty\}$ is isomorphic to the free group generated by two elements, this amounts to considering maps
$\rho:\pi_{1}\left(\mathbb{C}P^{1}\setminus\{0,1,\infty\}\right)\cong\langle a,b\rangle\rightarrow\textrm{GL}(3,\mathbb{C})$
under the constraint that the eigenvalues of $\rho(a),\rho(b)$ and $\rho(ab)$ are fixed: $\{\lambda_{1},\lambda_{2},\lambda_{3}\}$, $\{\mu_{1},\mu_{2},\mu_{3}\}$ and $\{\nu_{1},\nu_{2},\nu_{3}\}$ respectively (previously denoted by the vectors $\underline{c}_{P}$). Because of the $\textrm{PGL}(3,\mathbb{C})$ action, we have the freedom to choose $\rho(a)$ to be diagonal with elements $\{\lambda_{1},\lambda_{2},\lambda_{3}\}$. Having achieved this, there remains the action of the maximal torus (×$\left(\mathbb{C}^{\times}\right)^{2}$ of $\textrm{PGL}(3,\mathbb{C})$. The action of $(t_{1},t_{2})\in\left(\mathbb{C}^{\times}\right)^{2}$ on a matrix $B=[b_{ij}]$ is the standard conjugation
@@
$\mathcal{M}_{B}=\{A,B\in\textrm{GL}(3,\mathbb{C})|A=\textrm{Diag}[\lambda_{1},\lambda_{2},\lambda_{3}],B=[b_{i,j}],$
@@ 3.2 The trace coordinates
Notice that we used two different types of coordinates, but the two pictures coincide, as expected: for our special choice of parameters, the compactifying divisor of the Betti space is topologically the same as in the general description. According to Simpson [44], the Riemann Hilbert correspondence provides equivalence between vector bundles with integrable connection and representations of the fundamental group, via the monodromies of the connection around the punctures. Therefore our aim is to apply the trace coordinates to the monodromies of the connection, and investigate their asymptotic behaviour.
@@ 4 Asymptotic analysis of non-abelian Hodge correspondence
Consider a Higgs bundle $(\mathcal{E},\theta_{1})\in\mathcal{M}_{\textrm{Dol}}(\alpha)$ such that $H(\mathcal{E},\theta_{1})=q_{1}$ where $q_{1}$ is defined in (7). For any nonzero complex number $\tau$, the Higgs bundle (,¯$(\mathcal{E},\overline{\partial}_{E},\tau\theta_{1})$ is also stable with the same parabolic structure, therefore it gives another element of $\mathcal{M}_{\textrm{Dol}}(\alpha)$. For this Higgs bundle we clearly have $H(\mathcal{E},\tau\theta_{1})=\tau^{3}q_{1}=tq_{1}$, and the real Hitchin’s equation reads as
@@ 4.1.6 Hermitian metric on the spectral line bundle
Now, let us equip $\mathcal{L}\rightarrow\Sigma_{1}$ with a parabolic structure. For all the lifted points $\widetilde{P}\in\widetilde{D}$ on the spectral curve let us set the parabolic weights to be $\widetilde{\alpha}_{\widetilde{P}}=-1$.Then $\textrm{pdeg}\,\mathcal{L}=0$, because $\textrm{deg}\,\mathcal{L}=3$. By [46][2], there exists a unique metric $h_{\mathcal{L}}$ that
@@ 4.2.2 Flat unitary connection induced by $h_{\infty},\mathcal{E}$
[…]
@@ 4.3.4 Equivariance and scaling properties
5 Asymptotic analysis of the Riemann Hilbert correspondence
@@ 6.1 A topological reformulation of the Geometric P=W conjecture
$\sigma\colon S^{1}_{R}\to\mathcal{M}_{\textrm{Dol}}(\alpha)$
@@
Assume that $\operatorname{dim}_{\mathbb{R}}\mathcal{M}_{\textrm{Dol}}(\alpha)=4$. Assume that $S\circ\psi\circ\sigma$ induces an isomorphism
for $R\gg 1$. Then the diagram (1) is commutative up to homotopy.
@@
The argument of [39, Section 5] applies verbatim. It only depends on the condition that the moduli space is a $4$-manifold, not on the rank nor the type of singularities of the underlying Higgs bundles. ∎
Notice that in this proof we do not even need algebraicity (or even analyticity) of $\sigma$. Therefore, the proof of Theorem 1.1 reduces to the following.
@@
We devote the rest of this section to proving this assertion. Notice that the $S^{1}$ on the left-hand side does have a natural orientation coming from the complex structure of the Hitchin base, but the one on the right hand side does not have a preferred orientation.
@@ 6.2.5 Dominance between different trace coordinates within sectors
Figure 3: The complex plane $(a,b)$, sectors $S_{j}$, dominant coordinate $\lvert X_{R,\varphi}\rvert$ or $\lvert Y_{R,\varphi}\rvert$ depending on the sector.
@@ 6.2.6 Phases of the trace coordinates in terms of holonomy of the spectral line bundle
In Lemma 6.10, we determined the dominance properties of the absolute values of the trace coordinates along rays in any open sector $S_{j}$. We have found that $X/Z\to 0$ and $Y/Z\to 0$ as $R\to\infty$. These quantities constitute local affine coordinates near the nodal point of $C$. Here, we focus on the angles of these affine coordinates. Notice that a punctured neighborhood of the nodal point in the character variety admits a two-dimensional torus $T^{2}=S^{1}\times S^{1}$ as a deformation retract, whose $S^{1}$-factors correspond precisely to the angles of $X/Z$ and $Y/Z$. We find that the angles can be interpreted as the holonomy of $\mathcal{L}$ along some loops in $\Sigma$, depending on the given sector. Due to the $32\pi/$ 3 periodicity of the diagram of Figure 3, it is sufficient to deal with four consecutive sectors.
First, let us discuss the homology of $\widetilde{\Sigma}$ in some detail. Consider the straight line segment $[r,1-r]\subset\mathbb{R}$. Let us temporarily denote its lifts to $\Sigma$ by $f_{1},f_{2},f_{3}$, where $f_{j}$ lies on the sheet $\Sigma_{j}$ determined by the $Q_{j}$-eigenspace of $\theta$ (see (11)). On the other hand, the three lifts of the loop $\gamma_{0}$ cyclically connect the three sheets. For instance, one of these lifts connects $\Sigma_{1}$ to $\Sigma_{2}$. Denote this lift by $\tilde{\gamma}_{0}$. Moreover, let $\tilde{\tilde{\gamma}}_{0}$ denote the lift of $\gamma_{0}$ connecting the sheet $\Sigma_{2}$ to $\Sigma_{3}$. Similarly, one of the lifts $\tilde{\gamma}_{1}$ of $\gamma_{1}$ connects the sheet $\Sigma_{1}$ to $\Sigma_{2}$, and another one $\tilde{\tilde{\gamma}}_{1}$ connects $\Sigma_{2}$ to $\Sigma_{3}$. It follows that the paths $f_{1}*\tilde{\gamma}_{0}*(-f_{2})*(-\tilde{\gamma}_{1})$ and $f_{2}*\tilde{\tilde{\gamma}}_{0}*(-f_{3})*(-\tilde{\tilde{\gamma}}_{1})$ represent singular $1$-cycles on $\widetilde{\Sigma}$. Let us denote these cycles in order by $A,B$. See Figure 4.

### v2

@@ 2.3 The Hitchin fibration
According to Lemma 2.5, with the choices made in Section 2.2, the first two components vanish. The third characteristic coefficient of $\theta$ is $\operatorname{det}\theta\in H^{0}(\mathbb{C}P^{1},K(D)^{\otimes 3})$. Let us use the notation $L=K(D)$, with the natural projection from the total space of $L$, $p_{L}:\operatorname{Tot}L\rightarrow\mathbb{C}P^{1}$. Define $\zeta\frac{\operatorname{d}\!z}{z(z-1)}$ to be the canonical section of $p_{L}^{*}L$ over $p_{L}^{-1}(\mathbb{C})$, i.e. away from the infinity section. With this, the characteristic polynomial of $\theta$ is
@@
where $\operatorname{det}\theta$ may take its values is called the Hitchin base of $\mathcal{M}_{\operatorname{Dol}}(\alpha)$, and the map
[…]
@@
$\Sigma_{t}=\{(z,\zeta)|\zeta^{3}+tz(z-1)=0\}\subset\operatorname{Tot}L.$ (8)
Notice that $\Sigma_{t}$ has maximal ramification over $0,1$ and a smooth compactification, denoted by $\widetilde{\Sigma}_{t}$, in $\operatorname{Tot}L$ at $z=\infty$ with $(w,\zeta)=(0,0)$, where it also admits cyclic ramification. One can easily compute from Riemann–Hurwitz formula that its genus is equal to 1.
@@ 2.4 Ramification of the spectral curve and the Jacobian
$\operatorname{deg}\mathcal{L}=\operatorname{deg}\mathcal{E}+r(1-r)\left(1-g-\frac{n}{2}\right)=3,$
where $r=3$ is the rank, $n=3$ is the number of ramification points, and $g=0$ is the genus of $\mathbb{C}P^{1}$. Thus an element of the Dolbeault moduli space determines a spectral curve and a line bundle of degree 3 on it, and vice versa. So, the fiber of the Hitchin fibration over the point parameterized by a fixed $t\in\mathcal{H}$ is $\operatorname{Pic}^{3}(\widetilde{\Sigma}_{t})$, that is a 2-torus, namely a torsor over $\operatorname{Jac}(\widetilde{\Sigma}_{t})$. In particular, the Hitchin fibration is an elliptic fibration.
Let us discuss this correspondence between the Hitchin fibers and the Jacobian of the spectral curve a bit more in detail. Consider the following period lattice $\Lambda_{t}\subset H^{0,1}(\widetilde{\Sigma}_{t})\cong\mathbb{C}$, provided by the image $\operatorname{Im}(p^{0,1}\circ\iota)$, where
@@
are the coefficient inclusion on the first cohomology class, and the projection of harmonic forms to their antiholomorphic part respectively. There exists a $\mathbb{C}$-analytic isomorphism $\operatorname{Jac}(\widetilde{\Sigma}_{t})\cong H^{0,1}(\widetilde{\Sigma}_{t})/\Lambda_{t}$. Namely, any class in $H^{0,1}(\widetilde{\Sigma}_{t})/\Lambda_{t}$ can be represented by a $\mu\in\Omega^{0,1}(\widetilde{\Sigma}_{t})$, because of the abelian Hodge correspondence. Then the connection form $u=\mu-\overline{\mu}=2i\operatorname{Im}\mu\in\Omega^{1}(\widetilde{\Sigma}_{t})$ defines a flat $U(1)$-connection. For the line bundle $\mathcal{L}\in\operatorname{Jac}(\widetilde{\Sigma}_{t})$ given by $\mu$ and a $1$-cycle $X\in H_{1}(\widetilde{\Sigma}_{t},\mathbb{Z})$ we call
@@ 2.5 The Hitchin section
$\operatorname{res}_{P}(\tau\theta_{1})=\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&1&0\end{pmatrix}.$
@@ 3 Description of the Betti moduli space
In this chapter we consider the Betti space (or character variety), parameterizing the irreducible representations of the fundamental group of $\mathbb{C}P^{1}\setminus D$ in $\operatorname{GL}(3,\mathbb{C})$, up to the simultaneous conjugation by of $\operatorname{PGL}(3,\mathbb{C})$. We pick $z_{0}\in\mathbb{C}P^{1}\setminus\{0,1,\infty\}$ once and for all, and all occurrences of fundamental group will mean with base point $z_{0}$. Since the fundamental group of $\mathbb{C}P^{1}\setminus\{0,1,\infty\}$ is isomorphic to the free group generated by two elements, this amounts to considering maps
$\rho:\pi_{1}\left(\mathbb{C}P^{1}\setminus\{0,1,\infty\}\right)\cong\langle a,b\rangle\rightarrow\operatorname{GL}(3,\mathbb{C})$
under the constraint that the eigenvalues of $\rho(a),\rho(b)$ and $\rho(ab)$ are fixed: $\{\lambda_{1},\lambda_{2},\lambda_{3}\}$, $\{\mu_{1},\mu_{2},\mu_{3}\}$ and $\{\nu_{1},\nu_{2},\nu_{3}\}$ respectively (previously denoted by the vectors $\underline{c}_{P}$). Because of the $\operatorname{PGL}(3,\mathbb{C})$ action, we have the freedom to choose $\rho(a)$ to be diagonal with elements $\{\lambda_{1},\lambda_{2},\lambda_{3}\}$. Having achieved this, there remains the action of the maximal torus (×$\left(\mathbb{C}^{\times}\right)^{2}$ of $\operatorname{PGL}(3,\mathbb{C})$. The action of $(t_{1},t_{2})\in\left(\mathbb{C}^{\times}\right)^{2}$ on a matrix $B=[b_{ij}]$ is the standard conjugation
@@
$\mathcal{M}_{B}=\{A,B\in\operatorname{GL}(3,\mathbb{C})|A=\operatorname{Diag}[\lambda_{1},\lambda_{2},\lambda_{3}],B=[b_{i,j}],$
@@ 3.2 The trace coordinates
Notice that we used two different types of coordinates, but the two pictures coincide, as expected: for our special choice of parameters, the compactifying divisor of the Betti space is topologically the same as in the general description. According to Simpson [44], the Riemann–Hilbert correspondence provides equivalence between vector bundles with integrable connection and representations of the fundamental group, via the monodromies of the connection around the punctures. Therefore our aim is to apply the trace coordinates to the monodromies of the connection, and investigate their asymptotic behaviour.
@@ 4 Asymptotic analysis of non-abelian Hodge correspondence
Consider a Higgs bundle $(\mathcal{E},\theta_{1})\in\mathcal{M}_{\operatorname{Dol}}(\alpha)$ such that $H(\mathcal{E},\theta_{1})=q_{1}$ where $q_{1}$ is defined in (7). For any nonzero complex number $\tau$, the Higgs bundle (,¯$(\mathcal{E},\overline{\partial}_{E},\tau\theta_{1})$ is also stable with the same parabolic structure, therefore it gives another element of $\mathcal{M}_{\operatorname{Dol}}(\alpha)$. For this Higgs bundle we clearly have $H(\mathcal{E},\tau\theta_{1})=\tau^{3}q_{1}=tq_{1}$, and the real Hitchin’s equation reads as
@@ 4.1.6 Hermitian metric on the spectral line bundle
Now, let us equip $\mathcal{L}\rightarrow\Sigma_{1}$ with a parabolic structure. For all the lifted points $\widetilde{P}\in\widetilde{D}$ on the spectral curve let us set the parabolic weights to be $\widetilde{\alpha}_{\widetilde{P}}=-1$.Then $\operatorname{pdeg}\,\mathcal{L}=0$, because $\operatorname{deg}\,\mathcal{L}=3$. By [46][2], there exists a unique metric $h_{\mathcal{L}}$ that
@@ 4.2.2 Flat unitary connection induced by $h_{\infty},\mathcal{E}$
Then the flat $\operatorname{U(1)}^{\times 3}$-connection form on ${\left.\kern-1.2pt\mathcal{E}\vphantom{\big{|}}\right|_{U}}$, associated with $h_{\infty}$ is
@@ 4.3.4 Equivariance and scaling properties
5 Asymptotic analysis of the Riemann–Hilbert correspondence
@@ 6.1 A topological reformulation of the Geometric P=W conjecture
$\sigma\colon S^{1}_{R}\to\mathcal{M}_{\operatorname{Dol}}(\alpha)$
@@
Assume that $\operatorname{dim}_{\mathbb{R}}\mathcal{M}_{\operatorname{Dol}}(\alpha)=4$. Assume that for some $R\gg 1$ either
  1. 1.
    $S\circ\psi\circ\sigma$ induces an isomorphism
    $\mathbb{Z}\cong\pi_{1}(S^{1}_{R})\to\pi_{1}(\lvert\mathcal{D}\partial\mathcal{M}_{B}\rvert)\cong\mathbb{Z};$
    or
  2. 2.
    there exist regular values $z_{1},z_{2}\in S^{1}$ of $H$ and $S$ respectively such that $H^{-1}(Rz_{1})$ and $S^{-1}(z_{2})$ are both diffeomorphic to $T^{2}$ and that the inclusion map $H^{-1}(Rz_{1})\to\mathcal{M}_{\operatorname{Dol}}(\alpha)$ is homotopic to the inclusion map $S^{-1}(z_{2})\to\mathcal{M}_{\operatorname{B}}(\textbf{c})$, where the two ambient spaces are identified via $\psi$.
Then the diagram (1) is commutative up to homotopy.
Remark 6.2.
Part (2) generalizes to higher dimension via a framed cobordism argument, see [34, Theorem 5.4.1].
@@
Assuming the first condition, the argument of [39, Section 5] applies verbatim. It only depends on the condition that the moduli space is a $4$-manifold, not on the rank nor the type of singularities of the underlying Higgs bundles.
Assuming the second condition, the proof is again similar. Namely, the spaces $\mathcal{M}_{\operatorname{Dol}}(\alpha)\setminus H^{-1}(B_{R}(0))$ and $\mathcal{M}_{\operatorname{B}}(\textbf{c})\setminus\psi(H^{-1}(B_{R}(0)))$ deformation retract to their boundaries $Y_{\operatorname{Dol}}$ and $Y_{B}$ respectively. These spaces are compact $3$-manifolds, that are $T^{2}$-fibrations over $S^{1}$. The map $\psi$ is a homeomorphism between them, so we may refer to either of them by $Y$. Now, the set $[Y,S^{1}]$ of homotopy classes of maps $Y\to S^{1}$ is in bijection with $H^{1}(Y,\mathbb{Z})$ because $S^{1}=K(\mathbb{Z},1)$. On the other hand, by definition
$H^{1}(Y,\mathbb{Z})\cong\operatorname{Hom}(H_{1}(Y,\mathbb{Z}),\mathbb{Z}),$
so we get
$[Y,S^{1}]=\operatorname{Hom}(H_{1}(Y,\mathbb{Z}),\mathbb{Z})$
as sets. This bijection is given as follows: for a map $f\colon Y\to S^{1}$, take a regular value $z\in S^{1}$, and then we associate to $[f]$ the map
$[a]\in H_{1}(Y,\mathbb{Z})\mapsto[a]\cap[f^{-1}(z)]\in\mathbb{Z},$
where $\cap$ stands for algebraic intersection number. This finishes the proof, because the algebraic intersection number is homotopy-invariant. ∎
Notice that in this proof we do not even need algebraicity (or even analyticity) of $\sigma$. In view of Proposition 6.1, the proof of Theorem 1.1 reduces to the either of the following.
@@
Notice that the $S^{1}$ on the left-hand side does have a natural orientation coming from the complex structure of the Hitchin base, but the one on the right hand side does not have a preferred orientation.
Proposition 6.4.
There exist regular values $z_{1},z_{2}\in S^{1}$ of $H$ and $S$ respectively such that $H^{-1}(Rz_{1})$ and $S^{-1}(z_{2})$ are both diffeomorphic to $T^{2}$ and that the inclusion map $H^{-1}(Rz_{1})\to\mathcal{M}_{\operatorname{Dol}}(\alpha)$ is homotopic to the inclusion map $S^{-1}(z_{2})\to\mathcal{M}_{\operatorname{B}}(\textbf{c})$
We devote the rest of this section to proving these assertions. Namely, we will prove Proposition 6.3 conditional to an extra Assumption 2. We also show Proposition 6.4 unconditionally.
@@ 6.2.5 Dominance between different trace coordinates within sectors
Figure 3: The $\tau$ complex plane, sectors $S_{j}$, dominant coordinate $\lvert X_{R,\varphi}\rvert$ or $\lvert Y_{R,\varphi}\rvert$ depending on the sector.
@@
6.3 Proof of Proposition 6.4
6.3.1 Blow-up of the nodal point
Let us consider the blow up of the projective compactification of $\mathcal{M}_{\operatorname{B}}$ at the point $[0:0:1]$:
$\widetilde{\mathcal{M}_{\operatorname{B}}}\to\overline{\mathcal{M}_{\operatorname{B}}}.$
Locally around its nodal point $[0:0:1]$, the curve $C$ consists of two components, tangential to the $X$ and $Y$ axes respectively, in the local coordinate system $[X\colon Y\colon Z]$. Local affine coordinates near this point are given by $X/Z$ and $Y/Z$. Taking $\mathbb{C}P^{1}$ with homogeneous coordinates $[X^{\prime}\colon Y^{\prime}]$, the equation of the blown up surface reads as
$\frac{X}{Z}Y^{\prime}-\frac{Y}{Z}X^{\prime}=0.$
The inverse image of $(0,0)$ is the exceptional divisor, denoted by $C_{2}$. The proper transform of $C$ is denoted by $C_{1}$. Clearly, $C_{1}$ and $C_{2}$ have two intersection points. Denote the intersection point of $C_{2}$ with the proper transform of the $X$ axis by $P_{1}$, and the one with the $Y$ axis by $P_{2}$. The point $P_{1}$ is contained in the affine chart $U_{1}=\{Y^{\prime}\neq 0\}$. On $U_{1}$, the local affine coordinates are given by
$\frac{X}{Z},\quad\frac{X}{Y}.$
Similarly, $P_{2}$ is contained in the affine chart $U_{2}=\{X^{\prime}\neq 0\}$, and local affine coordinates are given by
$\frac{Y}{Z},\quad\frac{Y}{X}.$
The component $C_{i}$ gives rise to the vertex $V_{i}$ in the dual graph, and the two intersection points $P_{1},P_{2}$ correspond to edges $E_{1}$ and $E_{2}$ respectively. Then the dual boundary graph becomes as on Figure 4.
6.3.2 Simpson’s map
$S\colon\mathcal{M}_{\operatorname{B}}\setminus\psi(h^{-1}(B_{R}(0)))\to\lvert\mathcal{D}\partial\mathcal{M}_{B}\rvert$ (33)
as follows. We consider the compactification ~$\widetilde{\mathcal{M}}_{\operatorname{B}}$ of $\mathcal{M}_{\operatorname{B}}$ defined in Section 6.3.1. Next, we construct a cover
$\mathcal{M}_{\operatorname{B}}\setminus\psi(h^{-1}(B_{R}(0)))=N_{0}^{*}\cup N_{1}^{*}\cup N_{2}^{*}$ (34)
  1. 1.
    $N_{i}^{*}\cap N_{j}^{*}$ is non-empty and connected for every $0\leq i,j\leq 2$
  2. 2.
    $N_{0}^{*}\cap N_{1}^{*}\cap N_{2}^{*}=\emptyset$.
See Figure 7. Namely, let $N_{0}$ be a tubular neighborhood of $C_{1}$ in ~$\widetilde{\mathcal{M}}_{\operatorname{B}}$ and $N_{1},N_{2}$ be suitable neighborhoods of the intersection points $P_{1},P_{2}$ of $C_{1},C_{2}$ defined above. Let
$N_{i}^{*}=N_{i}\cap\mathcal{M}_{\operatorname{B}}=N_{i}\setminus(C_{1}\cup C_{2}).$
$\phi_{1}+\phi_{2}+\phi_{3}=1.$
Its edges are defined by the $\phi_{i}=0$ plane sections.
6.3.3 Homology basis
In this section, we will define a basis of $H_{1}(\widetilde{\Sigma},\mathbb{Z})$. Consider the straight line segment $[r,1-r]\subset\mathbb{R}$. Let us temporarily denote its lifts to $\Sigma$ by $f_{1},f_{2},f_{3}$, where $f_{j}$ lies on the sheet $\Sigma_{j}$ determined by the $Q_{j}$-eigenspace of $\theta$ (see (11)). On the other hand, the three lifts of the loop $\gamma_{0}$ cyclically connect the three sheets. For instance, one of these lifts connects $\Sigma_{1}$ to $\Sigma_{2}$. Denote this lift by $\tilde{\gamma}_{0}$. Moreover, let $\tilde{\tilde{\gamma}}_{0}$ denote the lift of $\gamma_{0}$ connecting the sheet $\Sigma_{2}$ to $\Sigma_{3}$. Similarly, one of the lifts $\tilde{\gamma}_{1}$ of $\gamma_{1}$ connects the sheet $\Sigma_{1}$ to $\Sigma_{2}$, and another one $\tilde{\tilde{\gamma}}_{1}$ connects $\Sigma_{2}$ to $\Sigma_{3}$. It follows that the paths $f_{1}*\tilde{\gamma}_{0}*(-f_{2})*(-\tilde{\gamma}_{1})$ and $f_{2}*\tilde{\tilde{\gamma}}_{0}*(-f_{3})*(-\tilde{\tilde{\gamma}}_{1})$ represent singular $1$-cycles on $\widetilde{\Sigma}$. Let us denote these cycles in order by $A,B$. See Figure 5.